a 78.2kg person is standing on a scale in a elevator. if the elevator is accelerating downwards at 0.307 m/s square what will be on the scale be ( in kilograms


1
Expert's answer
2021-03-18T14:02:28-0400
"m'=\frac{W'}{g}\\m'=\frac{m(g-a)}{g}\\m'=\frac{78.2(9.8-0.307)}{9.8}=75.75\ kg"